A number from 11 to 18 is drawn at random P (an odd number)
Found 2 solutions by stanbon, rfer:
Answer by stanbon(75887)

A number from 11 to 18 is drawn at random P (an odd number)
------
# of ways to succeed: 4
# of possible outcomes: 8
---
P(odd) = 4/8 = 1/2
====================
Cheers,
Stan H.

Answer by rfer(16322)